OpenStudy (anonymous):

How do you convert 2x + 3y = 9 into slope intercept form?

sam (.sam.):

Slope intercept form y=mx+c Try rearranging the equation 2x + 3y = 9

OpenStudy (anonymous):

so y=2 and then what do i do after that?

sam (.sam.):

No, rearrange it like this 2x + 3y = 9 3y = 9 - 2x 3y = -2x + 9 \[y=-\frac{2}{3}x+3\] Got it?

OpenStudy (anonymous):

yes, thank you for your help :)